OpenStudy (anonymous):

you have two containers. one is a cylinder with a height of 15 inches and a diameter of 10 inches. the other side is a cube with sides of 12 inches. what is the volume of the cylinder? what is the volume of the cube?

OpenStudy (anonymous):

volume of a cylendar = pi r^2 h. cube is length times width times depth

OpenStudy (anonymous):

so the volume of the cylinder would be 1177.5 inches^3

OpenStudy (anonymous):

i get 1178.09, lemme try it again though

OpenStudy (anonymous):

well for pie i just use the first 3 numbers.. 3.14

OpenStudy (anonymous):

ahh, yup that will do it

OpenStudy (anonymous):

and the cube would be 12*12=144

OpenStudy (anonymous):

that would only be the surface area of one side of the cube, its 12^3 height, width, and depth

OpenStudy (anonymous):

so 1728inches^3

OpenStudy (anonymous):

you got it

OpenStudy (anonymous):

one more question: how many gallons of water are needed to fill the cylinder? (hint: 1 gallon of water is .1337ft^3)

OpenStudy (anonymous):

if you know one gallon takes up .1337 ft cubed, and the cyl has 1177 ft cubed of space then some division will get you how many gallons per cyl

OpenStudy (anonymous):

and are you sure its .1337 per gallon? that seems kinda small considering a gallon of milk is roughly a cubic foot in real life

OpenStudy (anonymous):

that means you'd need 10 gallons to fill one cubic foot, just seems big to me but maybe im wrong?

OpenStudy (anonymous):

oooooh we are dealing in inches here. your .1137 ft needs to be converted to inches (easy, 12 inches per foot) to get you cubic inches or you can go the other way anf find out how many cubic feet the cyl has and then do your division. gotta make sure the scales are correct and then its simple division

OpenStudy (anonymous):

i cant convert worth crap i dont understand converting

OpenStudy (anonymous):

okay, if you have 12 inches per foot, theres 24 inches in 2 feet (12*2), theres 36 inches in 3 feet (12*3). so that also means you could take some random number of inches, lets pick 247 (totally random) and we know that 20.5833 feet is the same as 247 inches. (247/12). do you see how it works?

OpenStudy (anonymous):

multiplying and dividing are really just the same thing, one is forward and one is backwards. division tells you how many of A is in B. multiplying tells you if you had A number of groups and B number of items per group, what is the total number of items

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so i would do 1177.5*12

OpenStudy (anonymous):

almost, conversion of cubic numbers is a bit more trickey. you can either convert before finding the answer in ^3 or if you do it afterward you need to remember that your dealing with a volume in 3 dimensions. because think about the physical shape of a cubic foot (a box that is a ft by ft by ft) now think about 12 1 inch cubes (little bigger than 12 dice) not quite the same size as eachother. thats because your forgetting a whole dimensions

OpenStudy (anonymous):

however, if you know taht there are 1728 cubic inches per cubic foot (thats a true statement btw) then you can do your one to one conversions

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

does that make sense?

OpenStudy (anonymous):

kinda

OpenStudy (anonymous):

well to get that 1728 number all you are doing is saying, i have 12 units up, 12 units deep and 12 units wide.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so if you wanted to convert 4 cubic feet it would just be how many inches in one cubic foot times 4, aka 1728 times 4

OpenStudy (anonymous):

1177.5*4=4710

OpenStudy (anonymous):

sorry no, its not times 4 because that would be a 2X2X2 ft square. to find for a 4ft cubed would be 4*12=48 48^3=110592

OpenStudy (anonymous):

i meant to say 64 not 4 because it expands exponentially ie: 4^3 is what you multiply by (4^3=64)

OpenStudy (anonymous):

haha now i went and confused you more by typo-ing. im sorry about that. you have to remember that the number your multiplying by is cubed as well.

OpenStudy (anonymous):

im already confused

OpenStudy (anonymous):

its that 1728 number (how many cubic inches in a cubic foot) times how many cubic feet your finding ^3 example: 3 cubic feet has 1728*3^3 cubic inches in it. 24 cubic feet has 1728*24^3 cubic inches in it

OpenStudy (anonymous):

what i was saying before is you can figure it out by converting first if you want. so your 3 cubic feet has 3 feet in each direction. thats the same as 3*12=36 inches in each direction. now you can find cubic inces using those dimensions instead

OpenStudy (anonymous):

so 1177.5*3^3

OpenStudy (anonymous):

wait a second. your cyl you have in inch form and your gallons is in feet form. are you going to do your comparison in feet or inches because we need them both in the same format first

OpenStudy (anonymous):

ft

OpenStudy (anonymous):

so we need to divide 1177.5 in cubed by 12 inches cubed, that gets us feet

OpenStudy (anonymous):

oh! heres an idea that one of my teaches tried with us years ago, maybe that will help you with conversions, you can use it for any conversion you want

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

well apparently i have no idea how to use the equations thinger on this site, so thats not going to work. basically you just need to know how many units to convert and then write out equivalent fractions and then cross out matching units that cancel out.

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

http://www.purplemath.com/modules/units.htm

OpenStudy (anonymous):

i give up onthis one